# Work-From-Home and Privacy: What Do Workers Face and What are They Doing About it?

### v1

@@ 5. Conclusion
In this study, we surveyed 214 workers about their experiences with 14 privacy-invasive scenarios that can occur in a WFH setting. Our findings provide a better understanding of the privacy challenges that WFH workers face and how they address them, which can help in identifying ways to improve WFH experiences.
@@ 6. Acknowledgment
Eman Alashwali acknowledges the financial support of the Ibn Rushd Program at King Abdullah University of Science and Technology (KAUST). This work was funded in part by the Innovators Network Foundation. We thank Rachel Martini, Harish Balaji, Mahith Edula, and Zili Zhou for their contributions to the initial survey study during the Usable Privacy and Security course in 2023. We also thank the participants for their time and valuable input.
Appendix A Further Results
Table 5. Participants’ demographics.
N = 214
Age (years) No. %
From 18 to 24 17 7.9%
From 25 to 34 81 37.9%
From 35 to 44 67 31.3%
From 45 to 54 25 11.7%
From 55 to 64 21 9.8%
From 65 to 74 3 1.4%
75 or older 0 0.0%
Prefer not to answer 0 0.0%
Gender No. %
Male 116 54.2%
Female 94 43.9%
Non-binary 3 1.4%
Prefer to self describe: 0 0.0%
Prefer not to answer 1 0.5%
Race / Ethnicity No. %
White 154 72.0%
Black or African American 23 10.7%
American Indian or Alaska Native 7 3.3%
Asian 33 15.4%
Native Hawaiian or Pacific Islander 0 0.0%
Hispanic and/or Latino/Latina/Latinx 16 7.5%
Prefer to self describe: 3 1.4%
Prefer not to answer 3 1.4%
Annual Income No. %
Less than \$20,000 3 1.4%
\$20,000 to \$39,999 13 6.1%
\$40,000 to \$59,999 31 14.5%
\$60,000 to \$79,999 49 22.9%
\$80,000 to \$99,999 33 15.4%
\$100,000 to \$149,999 43 20.1%
\$150,000 or more 37 17.3%
Prefer not to answer 5 2.3%
Educatoin No. %
Doctoral degree 6 2.8%
Master’s degree 43 20.1%
Bachelor’s degree 108 50.5%
Associate’s degree 17 7.9%
High school diploma or GED 38 17.8%
Less than high school degree 0 0.0%
Other (please specify) 2 0.9%
CS/IS/IT/CE Background No. %
Yes 52 24.3%
No 162 75.7%
Cybersecurity Background No. %
Yes 16 7.5%
No 198 92.5%
Table 6. Participants’ work from home setup.
N = 214
Bedrooms in home No. %
1 bedroom 22 10.3%
2 bedrooms 54 25.2%
3 bedrooms 79 36.9%
4 bedrooms 42 19.6%
5 bedrooms 14 6.5%
More than 5 bedrooms 1 0.5%
Other (please specify) 2 0.9%
Living with (non-exclusive)
Spouse or partner 123 57.5%
Roommate(s) 15 7.0%
Child(ren) under the age of 18 68 31.8%
Child(ren) over the age of 18 17 7.9%
Parent(s) 36 16.8%
Sibling(s) 14 6.5%
No one 31 14.5%
Other (please specify) 6 2.8%
Duration living in the home No. %
Less than a year 16 7.5%
One year 22 10.3%
2 years 30 14.0%
3 years 21 9.8%
4 years 15 7.0%
5 years 6 2.8%
More than 5 years 102 47.7%
Other (please specify) 2 0.9%
Have a dedicated space for work No. %
Yes 170 79.4%
No 44 20.6%
Table 7. The privacy invasive scenarios ordered by categories. Columns 2–4 respectively list the number of participants who experienced the scenario and felt very uncomfortable; somewhat uncomfortable; and comfortable. The last column shows the percentage of participants who experienced the scenario and felt uncomfortable with respect to the number of participants who experienced that scenario (ratio of discomfort). Refer to Table 1 for description of the scenarios codes.
N = 214
Very uncomf. Somewhat uncomf. Comf.
Audio Scenarios
voice_you 17 44 39 114 61 100 61.0%
voice_adult 12 35 50 117 47 97 48.5%
voice_child 5 20 33 156 25 58 43.1%
sound_object 10 45 85 74 55 140 39.3%
sound_pet 9 42 71 92 51 122 41.8%
Video Scenarios
video_you 8 23 23 160 31 54 57.4%
video_adult 5 15 33 161 20 53 37.7%
video_child 6 8 25 175 14 39 35.9%
video_object 1 21 92 100 22 114 19.3%
video_pet 0 11 69 134 11 80 13.8%
Data Scenarios
computer_data 6 25 46 137 31 77 40.3%
browsing_data 4 30 35 145 34 69 49.3%
Autonomy Scenarios
cannot_stop_camera 9 33 15 157 42 57 73.7%
cannot_stop_mic 11 26 19 158 37 56 66.1%
N = 214
Scenario cat. Experienced Uncomf.
Audio 187 107
Video 138 55
Data 91 54
Autonomy 72 47
N = 140
Scenario No. %
Sound_pet 28 20.0%
Voice_you 19 13.6%
Voice_adult 14 10.0%
Sound_object 13 9.3%
Cant_stop_camera 12 8.6%
Video_you 11 7.9%
Computer_data 9 6.4%
Browsing_data 7 5.0%
Voice_child 7 5.0%
Cant_stop_mic 7 5.0%
Video_adult 5 3.6%
Video_child 4 2.9%
Video_object 2 1.4%
Video_pet 2 1.4%
Table 10. Characteristics of scenarios that caused discomfort.
N = 140
Discomfort escalated to cause harm No. %
It caused a large amount of harm 1 0.7%
It caused a moderate amount of harm 0 0.0%
It caused a small amount of harm 35 25.0%
It did not cause any harm 104 74.3%
Type of harm (non-exclusive; N = 36) No. %
Financial harm 1 2.8%
Physical harm 1 2.8%
Psychological harm 34 94.4%
Other (please specify) 1 2.8%
Frequency of event No. %
Once 71 50.7%
A few times 66 47.1%
Repeatedly 3 2.1%
I cannot remember 0 0.0%
Recording of event No. %
Yes, video and voice 25 17.9%
Yes, video only 4 2.9%
Yes, voice only 15 10.7%
No, not recorded 88 62.9%
I cannot remember 8 5.7%
Where in the house No. %
Office room 59 42.1%
Living room / recreation area 26 18.6%
Bedroom 30 21.4%
Kitchen 7 5.0%
Dining room 8 5.7%
Garage 0 0.0%
Bathroom 1 0.7%
Other (please specify) 9 6.4%
Device shared (non-exclusive) No. %
Microphone 115 82.1%
Camera 82 58.6%
Screen 28 20.0%
I cannot remember 1 0.7%
Other (please specify) 2 1.4%
Appendix B Surveys
In this section we include the surveys we used in the study. The subsections and text between [square brackets] were not shown to participants. We include them here for clarity. Appendix B.1 lists the recruitment materials. Appendix B.2 lists the screening survey, and Appendix B.3 lists the main study’s survey.
B.1. Recruitment Materials
@@
Description: Our research team at Anonymous Institution is searching for people to participate in a survey about their Work From Home Experiences. Participants should be working from home to participate.
@@
B.2. Screening Survey
B.2.1. Consent
@@ B.2.2. Introduction
[Q.18 is displayed if screening criteria are met:
@@
AND if Q.13 answer is (“Video conferencing (e.g. Zoom, Google Meet, Microsoft Teams, Skype, etc.)” AND (“Mobile smartphone” OR “Tablet” OR “Laptop” OR “Personal Computer (PC)”))]
@@
B.3. Main Survey
B.3.1. Introduction
@@ B.3.3. WFH Scenarios
[Q.6, Q.9, Q.12, and Q.15 are presented as matrices in randomized-order (the matrices), one matrix per page.]
@@
[Q.6, Q.9, Q.12, and Q.15 lists represents the matrices rows]
@@
[If one or more of Q.6 answers were (“Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”)]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If all Q.6 answers were “Never experienced it”]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If one or more of Q.9 answers were (“Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”)]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If all Q.9 answers were “Never experienced it”]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If one or more of Q.12 answers were (“Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”)]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If all Q.12 answers were “Never experienced it”]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If one or more of Q.15 answers were (“Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”)]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[If all Q.15 answers were “Never experienced it”]
@@
  • Yes (what are they?) Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@ B.3.4. Detailed Scenario
[Participants were shown a list of scenarios that they reported they made them feel “very uncomfortable” or “somewhat uncomfortable”]
@@
[Q.31 - Q.32 are shown if the participants did not select any scenario as “very uncomfortable” or “somewhat uncomfortable” in Q.6, Q.9, Q.12, and Q.15]
@@ B.3.5. Demographics
Q.37: What is your approximate annual household income? Please answer based on your entire current annual household’s income, before taxes.

### v3

@@ 6. Conclusion
In this study, we surveyed 214 workers about their experiences with 14 privacy-invasive scenarios that can occur in a WFH setting. Our findings provide a better understanding of the privacy challenges that WFH workers face and how they address them, which can help identify ways to improve WFH experiences.
7. Competing Interests
No competing interest to declare.
8. Author Contributions Statement
E. A. (Conceptualization, Data curation, Formal analysis, Funding acquisition, Investigation, Methodology, Visualization, Writing – original draft, Writing – review & editing); J. P. (Conceptualization, Data curation, Methodology, Writing – original draft); M. L. (Conceptualization, Data curation, Formal analysis, Methodology, Writing – original draft); L. F. c. (Conceptualization, Funding acquisition, Methodology, Writing – review & editing)
@@ 9. Acknowledgment
We thank Prof. Marc Dacier from KAUST for feedback on earlier versions of some sections of this paper. We also thank Harish Balaji, Mahith Edula, Rachel Martini, and Zili Zhou for their contributions to the initial survey study during the Usable Privacy and Security course at CMU. Finally, we thank the participants for their time and valuable input.
10. Funding
E. A. acknowledges the financial support of the Ibn Rushd Program at King Abdullah University of Science and Technology (KAUST). This work was supported in part by the Innovators Network Foundation (L. F. C.).
Appendix A Appendix A - Surveys
In this section we include the surveys we used in the study. The subsections and text between [square brackets] were not shown to participants. We include them here for clarity. Appendix A.1 lists the recruitment materials. Appendix A.2 lists the screening survey, and Appendix A.3 lists the main study’s survey.
A.1. RECRUITMENT MATERIALS
@@
Description: Our research team at Carnegie Mellon University is searching for people to participate in a survey about their Work From Home Experiences. Participants should be working from home to participate.
@@
A.2. SCREENING SURVEY
A.2.1. CONSENT
@@ A.2.2. INTRODUCTION
  • Other (please specify): [open-text entry]
@@
[Q.18 is displayed if the following screening criteria are met:
@@
AND if Q.13 answer is“Mobile smartphone” OR “Tablet” OR “Laptop” OR “Personal Computer (PC)”
AND if Q.14 answer is “Video conferencing (e.g. Zoom, Google Meet, Microsoft Teams, Skype, etc.)” OR (“Phone” OR “Email messaging” OR “Instant messaging applications (Slack, WhatsApp, etc.)” OR “Other (please specify)”)]
@@
A.3. MAIN SURVEY
A.3.1. INTRODUCTION
@@ A.3.3. WFH SCENARIOS
[Q.6, Q.9, Q.12, and Q.15 are displayed as matrices in randomized-order (the matrices), one matrix per page.]
@@
[Q.6, Q.9, Q.12, and Q.15 scenario lists represents the matrices rows]
@@
[Q.7 is displayed if one or more of Q.6 answers were “Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.8 is displayed if all Q.6 answers were “Never experienced it”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.10 is displayed if one or more of Q.9 answers were “Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.11 is displayed if all Q.9 answers were “Never experienced it”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.13 is displayed if one or more of Q.12 answers were “Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.14 is displayed if all Q.12 answers were “Never experienced it”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.16 is displayed if one or more of Q.15 answers were “Experienced it and felt very uncomfortable” OR “Experienced it and felt somewhat uncomfortable” OR “Experienced it and felt comfortable”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@
[Q.17 is displayed if all Q.15 answers were “Never experienced it”]
@@
  • Yes (what are they?)
    Please describe any protective measures you took, including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.
@@ A.3.4. DETAILED SCENARIO
[A list of the scenarios that the participant reported that they made them feel “very uncomfortable” or “somewhat uncomfortable” in Q.6, Q.9, Q.12, and Q.15 is displayed here]
@@
[Q.31 - Q.32 are displayed if the participants did not select any scenario as “very uncomfortable” or “somewhat uncomfortable” in Q.6, Q.9, Q.12, and Q.15]
@@ A.3.5. Demographics
Q.37: What is your approximate annual household income?
Please answer based on your entire current annual household’s income, before taxes.
@@
Appendix B Appendix B - Further Results
B.1. Screening Survey Results
Table 5 and Table 6 list the screening survey results for our participants.
Table 5. Participants’ work context.
N = 214
Employment type No. %
Employee (full-time) 175 81.8%
Employee (part-time) 19 8.9%
Self-employed / Freelancer / Business-owner 20 9.3%
Work sector No. %
Pre-university Education 3 1.4%
University Education 11 5.1%
Health 23 10.7%
Information and Communication Technology 48 22.4%
Financial 29 13.6%
Industrial 7 3.3%
Agricultural 0 0.0%
Sales and retail 26 12.1%
Petrochemical 0 0.0%
Other (please specify) 67 31.3%
Organization type No. %
Privately-held organization 116 54.2%
Publicly-traded organization 41 19.2%
Government organization 25 11.7%
Educational organization 14 6.5%
Not-for-profit organization 17 7.9%
Other (please specify) 1 0.5%
Duration working in current job No. %
Less than a year 16 7.5%
One year 23 10.7%
2 years 42 19.6%
3 years 20 9.3%
4 years 14 6.5%
5 years 18 8.4%
More than 5 years 81 37.9%
Other (please specify) 0 0.0%
Avg. normal work days/week No. %
One day per week 1 0.5%
2 days per week 3 1.4%
3 days per week 7 3.3%
4 days per week 11 5.1%
5 days per week 177 82.7%
6 days per week 12 5.6%
7 days per week 3 1.4%
Other (please specify) 0 0.0%
Avg. WFH days/week No. %
One day per week 18 8.4%
2 days per week 28 13.1%
3 days per week 33 15.4%
4 days per week 31 14.5%
5 days per week 95 44.4%
6 days per week 7 3.3%
7 days per week 2 0.9%
Other (please specify) 0 0.0%
Duration been WFH No. %
Less than 3 months 0 0.0%
From 3 to 11 months 21 9.8%
One year 23 10.7%
2 years 45 21.0%
3 years 66 30.8%
4 years 27 12.6%
5 years 9 4.2%
More than 5 years 23 10.7%
Other (please specify) 0 0.0%
Table 6. Participants’ conference tools usage.
N = 214
Frequency comm. via conference tools No. %
At least once every work from home day 132 61.7%
At least once every few work from home days 43 20.1%
At least once every work from home week 28 13.1%
At least once every few work from home weeks 11 5.1%
At least once every work from home month 0 0.0%
Less than once every work from home month 0 0.0%
Other (please specify) 0 0.0%
Avg. time spent comm via conference tools No. %
Less than one hour per work from home day 51 23.8%
At least one hour per work from home day 82 38.3%
At least 2 hours per work from home day 46 21.5%
At least 3 hours per work from home day 19 8.9%
At least 4 hours per work from home day 6 2.8%
More than 4 hours per work from home day 10 4.7%
Other (please specify) 0 0.0%
Devices used in WFH (non-exclusive) No. %
Mobile smartphone 151 70.6%
Mobile traditional (non-smart) phone 0 0.0%
Landline phone 13 6.1%
Tablet 28 13.1%
Laptop 194 90.7%
Personal Computer (PC) 74 34.6%
Other (please specify) 0 0.0%
Techs used in WFH (non-exclusive) No. %
Video conferencing (e.g. Zoom, Google Meet, Microsoft Teams, Skype, etc.) 214 100.0%
Phone (audio only) 124 57.9%
Email messaging 194 90.7%
Instant messaging applications (Slack, WhatsApp, etc.) 141 65.9%
Mobile Short Messaging Service (SMS) messaging 79 36.9%
Other (please specify) 3 1.4%
Frequency share microphone No. %
All of my meetings 75 35.0%
Most of my meetings 74 34.6%
About half of my meetings 30 14.0%
A few of my meetings 31 14.5%
None of my meetings 4 1.9%
Frequency share camera No. %
All of my meetings 46 21.5%
Most of my meetings 56 26.2%
About half of my meetings 32 15.0%
A few of my meetings 62 29.0%
None of my meetings 18 8.4%
Frequency share screen No. %
All of my meetings 8 3.7%
Most of my meetings 20 9.3%
About half of my meetings 41 19.2%
A few of my meetings 121 56.5%
None of my meetings 24 11.2%
B.2. Main Survey Results
Table 7 and Table 8 list demographic details and WFH setting for our participants.
Table 7. Participants’ demographics.
N = 214
Employer productivity monitoring (non-exclusive) No. %
Time trackers 41 19.2%
Activity trackers 39 18.2%
Task trackers 30 14.0%
Video monitoring 11 5.1%
Audio monitoring 10 4.7%
I am not aware of any productivity trackers used by my employers 143 66.8%
Other (*please specify) 6 2.8%
Age (years) No. %
From 18 to 24 17 7.9%
From 25 to 34 81 37.9%
From 35 to 44 67 31.3%
From 45 to 54 25 11.7%
From 55 to 64 21 9.8%
From 65 to 74 3 1.4%
75 or older 0 0.0%
Prefer not to answer 0 0.0%
Gender No. %
Male 116 54.2%
Female 94 43.9%
Non-binary 3 1.4%
Prefer to self describe: 0 0.0%
Prefer not to answer 1 0.5%
Race / Ethnicity (non-exclusive) No. %
White 154 72.0%
Black or African American 23 10.7%
American Indian or Alaska Native 7 3.3%
Asian 33 15.4%
Native Hawaiian or Pacific Islander 0 0.0%
Hispanic and/or Latino/Latina/Latinx 16 7.5%
Prefer to self describe: 3 1.4%
Prefer not to answer 3 1.4%
Annual Income No. %
Less than \$20,000 3 1.4%
\$20,000 to \$39,999 13 6.1%
\$40,000 to \$59,999 31 14.5%
\$60,000 to \$79,999 49 22.9%
\$80,000 to \$99,999 33 15.4%
\$100,000 to \$149,999 43 20.1%
\$150,000 or more 37 17.3%
Prefer not to answer 5 2.3%
Education No. %
Doctoral degree 6 2.8%
Master’s degree 43 20.1%
Bachelor’s degree 108 50.5%
Associate’s degree 17 7.9%
High school diploma or GED 38 17.8%
Less than high school degree 0 0.0%
Other (please specify) 2 0.9%
CS/IS/IT/CE Background No. %
Yes 52 24.3%
No 162 75.7%
Cybersecurity Background No. %
Yes 16 7.5%
No 198 92.5%
Table 8. Participants’ WFH setting.
N = 214
Bedrooms in home No. %
1 bedroom 22 10.3%
2 bedrooms 54 25.2%
3 bedrooms 79 36.9%
4 bedrooms 42 19.6%
5 bedrooms 14 6.5%
More than 5 bedrooms 1 0.5%
Other (please specify) 2 0.9%
Living with (non-exclusive) No. %
Spouse or partner 123 57.5%
Roommate(s) 15 7.0%
Child(ren) under the age of 18 68 31.8%
Child(ren) over the age of 18 17 7.9%
Parent(s) 36 16.8%
Sibling(s) 14 6.5%
No one 31 14.5%
Other (please specify) 6 2.8%
Duration living in the home No. %
Less than a year 16 7.5%
One year 22 10.3%
2 years 30 14.0%
3 years 21 9.8%
4 years 15 7.0%
5 years 6 2.8%
More than 5 years 102 47.7%
Other (please specify) 2 0.9%
Have a dedicated space for work No. %
Yes 170 79.4%
No 44 20.6%
Table 9. The privacy invasive scenarios ordered by categories. Columns 2–4 respectively list the number of participants who experienced the scenario and felt very uncomfortable; somewhat uncomfortable; and comfortable. The last column shows the percentage of participants who experienced the scenario and felt uncomfortable with respect to the number of participants who experienced that scenario (ratio of discomfort). Refer to Table 1 in the main paper for description of the scenarios codes.
N = 214
Very uncomf. Somewhat uncomf. Comf.
Audio Scenarios
voice_you 17 44 39 114 61 100 61.0%
voice_adult 12 35 50 117 47 97 48.5%
voice_child 5 20 33 156 25 58 43.1%
sound_object 10 45 85 74 55 140 39.3%
sound_pet 9 42 71 92 51 122 41.8%
Video Scenarios
video_you 8 23 23 160 31 54 57.4%
video_adult 5 15 33 161 20 53 37.7%
video_child 6 8 25 175 14 39 35.9%
video_object 1 21 92 100 22 114 19.3%
video_pet 0 11 69 134 11 80 13.8%
Data Scenarios
computer_data 6 25 46 137 31 77 40.3%
browsing_data 4 30 35 145 34 69 49.3%
Autonomy Scenarios
cannot_stop_camera 9 33 15 157 42 57 73.7%
cannot_stop_mic 11 26 19 158 37 56 66.1%
N = 214
Scenario cat. Experienced Uncomf.
Audio 187 107
Video 138 55
Data 91 54
Autonomy 72 47
N = 140
Scenario No. %
sound_pet 28 20.0%
voice_you 19 13.6%
voice_adult 14 10.0%
sound_object 13 9.3%
cant_stop_camera 12 8.6%
video_you 11 7.9%
computer_data 9 6.4%
browsing_data 7 5.0%
voice_child 7 5.0%
cant_stop_mic 7 5.0%
video_adult 5 3.6%
video_child 4 2.9%
video_object 2 1.4%
video_pet 2 1.4%
Table 12. Characteristics of scenarios that caused discomfort.
N = 140
Discomfort escalated to cause harm No. %
It caused a large amount of harm 1 0.7%
It caused a moderate amount of harm 0 0.0%
It caused a small amount of harm 35 25.0%
It did not cause any harm 104 74.3%
Type of harm (non-exclusive; N = 36) No. %
Financial harm 1 2.8%
Physical harm 1 2.8%
Psychological harm 34 94.4%
Other (please specify) 1 2.8%
Frequency of event No. %
Once 71 50.7%
A few times 66 47.1%
Repeatedly 3 2.1%
I cannot remember 0 0.0%
Recording of event No. %
Yes, video and voice 25 17.9%
Yes, video only 4 2.9%
Yes, voice only 15 10.7%
No, not recorded 88 62.9%
I cannot remember 8 5.7%
Where in the house No. %
Office room 59 42.1%
Living room / recreation area 26 18.6%
Bedroom 30 21.4%
Kitchen 7 5.0%
Dining room 8 5.7%
Garage 0 0.0%
Bathroom 1 0.7%
Other (please specify) 9 6.4%
Device shared (non-exclusive) No. %
Microphone 115 82.1%
Camera 82 58.6%
Screen 28 20.0%
I cannot remember 1 0.7%
Other (please specify) 2 1.4%